write the point-slope form of the equation of the line that passes through the two points


(4, -2), (-9, 5)
Thanks for help

Answer by checkley71(8403)

FIRST YOU NEED TO FIND THE SLOPE (Y2-Y1)/(X2-X1)
(5+2)/(-9-4)=7/-13 OR -7/13
Y=mX+b.
NOW SUPSTITUTE ONE SET OF X,Y POINTS & THE SLOPE (m) & SOLVE FOR (b) THE Y INTERCEPT.
-2=-7/13*4+b
-2=-28/13+b
b=-2+28/13
b=(-26+28)/13
b=2/13 ANSER FOR THE Y INTERCEPT.
THUS WE HAVE THE FOLLOWING LINE EQUATION:
Y=-7X/13+2/13